Suppose a perfectly competitive firm produces 2 outputs. The firm’s price for the first output is P1 and the price for the second output is P2. The cost function is given by:

C(Q1, Q2) = 2Q12 + 2Q22

a) Give the profit function for the firm.

b) Find the FOC’s for profit maximization and interpret them economically.

c) Find the SOC’s.
Answer #1

a) IT = = Ploit P202 - c(01 02) Piol + P2O2- 201 - 202 - Pi and P2 firm is which firm is are constant because porteatty compeo soci 020 220 oo? 021 201202 = Now the ) - ( 220 2012 oz or) 00% E 1-4) (-4) - 1022__ = 16-0 = 16 > since 02 siuce la posta